# Towards AI-Architecture Liberty: A Comprehensive Survey on Designing and Collaborating Virtual Architecture by Deep Learning in the Metaverse

### v3

@@ Abstract.
3D shape generation techniques leveraging deep learning have garnered significant interest from both the computer vision and architectural design communities, promising to enrich the content of the future metaverse. However, research on virtual architectural design remains limited, particularly regarding human-AI collaboration and deep learning-assisted design. We first illuminate the principles, generation techniques, and current literature of virtual architecture, focusing on challenges such as datasets, multimodality, design intuition, and generative frameworks. In our survey, we reviewed 187 related articles (80.7% of articles published between 2018 and 2022) covering architectural research, virtual environments, and technical approaches. This survey investigates the latest approaches to 3D object generation with deep generative models (DGMs) and summarizes four characteristics of deep-learning generation approaches for virtual architecture. According to our analysis of the survey, we expound on four research agendas, including agency, communication, user consideration, and integrating tools, and highlight three important enablers of ubiquitous interaction with immersive systems in deep learning-assisted architectural generation. Our work contributes to fostering understanding between designers and deep learning techniques, broadening access to human-AI collaboration. We advocate for interdisciplinary efforts to address this timely research topic, facilitating content designing and generation in the metaverse.
Deep Learning, virtual environment, architectural design, computational architecture, 3D shape generation, 3D-aware image synthesis, human-computer interaction, metaverse, AIGC
@@ 1. Introduction
In recent decades, the exploration of architectural space has shifted from reinforced concrete to digital architecture and further to information frameworks, with virtualization extending beyond the physical layer (Carpo, 2017; Bava, 2020). Architecture is no longer perceived as static, permanent objects but as integral components of a data network and evolving communication between different architectural systems (Claypool, 2019; Schroeder et al., 2001). Digital innovations and technological advancements associated with spline, pixels, voxels, and bits have redefined architectural forms called digital architecture (Jovanovic, 2022). Given this sense, scholars and practitioners have called AI architecture a critical sub-genre of architectural discourse (Stojanovski et al., 2022), which takes neural networks in deep learning concerning social issues. Moreover, the burgeoning metaverse has expanded the boundaries of production and design in digital fabrication buildings, contributing to the development of virtual architecture from various aspects, such as bio-signal feedback (Ding et al., 2022a; Pei et al., 2020; Homolja et al., 2020; Pei et al., 2021), physical-virtual merging (Cheng et al., 2019; Lee et al., 2021a), design assistance (Lin and Lo, 2021), design generation (Jang and Kim, 2024), planning principle (Ingram et al., 1996), and public participation (Kim et al., 2022; Kwiecinski et al., 2017). In this context, deep learning facilitates the generation and design of diverse virtual architectures with interactive features, depending on ubiquitous intelligent environments in the metaverse. It simulates real-world scenarios by generating and utilizing 3D building datasets, aligning design objectives with human needs (Özerol and Selçuk, 2023).
@@
Figure 1. Virtual architecture provides rich use affordances, benefiting various aspects, including users, communities, and organizations. (a) Hub Rooms in VRChat; (b) Simulated city scene with various buildings; (c) Interior of individual house in a virtual world game, Second Life; (d) Saint Motel Fan Meet on Mozzila hub; (e) Sci-fi nightclub; (f) Center building in pride month event held on Dencentraland; (g) Realism Sotheby’s virtual gallery; (h) The interactive digital experience in a virtual building for the brand Jose Cuervo designed by Rojkind Arquitectos; (i) The office building of Vice Media Group was designed in metaverse by BIG Architect; (j) ”NFTism” is a virtual NFT gallery developed collaboratively by Zaha Hadid Architects (ZHA) and Journee. It is a virtual NFT gallery named ”NFTism.” The architectural design focuses on user experience, social interaction, and ”dramaturgical” compositions, supporting MMO (massively multiplayer online) technology and integrating audio-video interaction [^2] A report by Archdaily:https://www.archdaily.com/972886/zaha-hadid-architects-presents-virtual-gallery-exploring-architecture-nfts-and-the-metaverse; (k) The virtual city designed and developed by ZHA, filled with fluid style virtual architecture; (l) Each piece of land and each item in the virtual land as a non-fungible token could trade in the virtual world of Decentraland; (m) A mapping showing ownership of non-fungible token land in Sandbox; (n) A user-friendly tool for editing and creating buildings and land in social VR, Mozzila hub.
@@
Integrating virtual buildings with deep generative models through innovative designer-AI collaboration brings significant benefits to the metaverse across various dimensions. Firstly, generated virtual buildings contribute to efficient and massive production, expanding metaverse scenarios with a variety of content (Jang and Kim, 2024). Secondly, generated virtual buildings prioritize user experiences by considering spatial, social, and physical aspects within the context of human-building interaction (HBI), such as evolving spatial forms and adaptive architectural elements (Alavi et al., 2019). Thirdly, generated virtual buildings offer social sustainability and economic benefits by embedding computational frameworks or workflows automatically (Jang and Kim, 2024), aligning with development needs in the metaverse. These benefits include time and resource savings, as well as fostering continuous renewal and evolutionary capabilities. Additionally, generated virtual architecture with deep learning conveniently provides access for both experts and inexperienced users in content production through AI-human collaboration, contributing to the content ecosystem in the metaverse. One concrete example is that the user could create a house or edit land using the user-friendly tool in social VR (Fig. 2n).
@@
Deep learning’s powerful capability enables the fusion of virtual and physical realities, facilitating diverse lifestyles and a wide range of user, community, and organizational uses (as shown in Fig. 2). Virtual buildings could enhance user presence by immersing themselves in vivid, realistic scenes of the metaverse (Jang and Kim, 2024) (Fig.  2a-c). Virtual buildings also serve as expansive communication interfaces and containers, fostering social activities and goals in immersive environments (Schroeder et al., 2001). For instance, users require virtual buildings to hold various activities, events, and communities, such as fan meets (Fig. 2d), club parties (Fig. 2e), and Pride Month parades (Fig. 2f). Organizations also require virtual buildings to maintain urban business premises, including companies, cultural institutions, and charity organizations (Fig. 2g-i). Notably, renowned architectural firms have begun designing virtual cities and buildings in virtual worlds (Fig. 2h-k). Furthermore, the interactive nature of virtual architecture facilitates diverse social systems. It delivers sophisticated information regarding social interactive elements in multiple scenarios, as demonstrated by ZAHA’s advanced virtual architectural projects incorporating interactive technology and architectural design (Fig. 2h-j). Lastly, virtual architecture enables trading, giveaways, and auctions of virtual land as non-fungible tokens (Fig. 2l-m), enriching the metaverse ecosystem.
@@
The initial applications of emerging technology often face access issues, and knowledge from one field may not transfer to another (Davies et al., 2018). In architecture, design research in architecture predominantly focuses on generative images like floor plans (Chaillou, 2022, 2020), exterior facades (Mohammad et al., 2019), and stylized images (del Campo et al., 2019; Del Campo, 2021), leveraging neural networks like style GAN and pix2pix. However, the potential of deep learning for 3D shape generation in virtual building applications remains largely overlooked. Despite proposals for computational and algorithmic approaches (e.g., (Jang and Kim, 2024; Lab, 2023)), design dimensions of virtual architecture have yet to be systematically explored. This gap impedes the advancement of virtual buildings in the metaverse. Our survey aims to bridge this gap by investigating deep-learning-based 3D shape generation methods for virtual architecture. We emphasize the importance of considering human-AI collaboration perspectives, particularly inclusivity, to cater to non-tech-savvy architects and layperson users seeking to create virtual buildings in the metaverse.
@@ 1.2. Methodology and Related Articles
Figure 3. (a) This survey investigates this intersection area. (b) A profile of the number of works cited in this paper in different categories and years. The survey’s scope and profile of related articles: a – architectural studies on DGMs; c – computer vision studies; v – those works on rules in VWs.
As abovementioned problems, the intersection of 3D generation techniques and virtual architecture is still nearly blank. Therefore, we anchored the three fields to conduct this survey to complement the key insight with each other: 3D shape generation techniques, DL-assisted architectural design, and the design considerations in a VWs in terms of HCI (see Fig. 2(a)). This article conducts data collection by utilizing the Preferred Reporting Items for Systematic Reviews and Meta-Analyses (PRISMA) framework (Page et al., 2021).
Identification. We reviewed a sample of 187 articles and primarily focused on works published between 2018 and 2022 (five years, 80.7%) as follows: 2023 or later: 4 (8.7%), 2022: 33 (17.6%), 2021: 39 (20.9%), 2020: 36 (19.3%), 2019: 30 (16%), 2018: 33 (5.8%), before 2018: 32 (17.1%) from those three fields (see Fig.  2(b)). We found the articles primarily through publication databases such as ACM Digital Library, IEEE Xplore, ScienceDirect, Springer Link, and CuminCAD. These databases ensure a wide range of computer graphics, virtual environments, and architectural design scope. We determined our search query, including the primary and second terms as searching keywords, after discussion with inner researchers through multiple sessions of iteration and refinements. These keywords were inspired by previous review papers. The primary keywords in the search query are straightforward reflecting our survey topic, as follows:
@@
    Articles discussing architectural design solely within the realm of reality (e.g., natural environment, BIM, construction, floor plans) were excluded.
@@
    Articles whose primary focus was not on 3D design generation (e.g., analysis of design cognition or limited 2D generative drawing without 3D transition approaches) were excluded.
@@
    Articles primarily discussing generative approaches other than deep learning for architecture (e.g., reinforcement learning, genetic algorithms) were excluded. Only articles discussing deep generative models for building generation were considered.
@@
    Articles discussing wide-ranging architecture design topics (e.g., the architecture of algorithms) were excluded.
@@
    Articles related to 3D generation focusing on the classification and retrieval of computer vision approaches were excluded.
When the keywords and abstracts do not appear as the key information or elements in our investigating scope, we read the whole publication to check whether it is included or not. After the screening, we got 147 articles and 19 CV research published on arXiv to review, i.e., 166 authoritative articles. Additionally, online resources were directly searched through the Google search engine, and we mainly concluded 19 articles and 2 relevant architectural projects from the perspective of architecture design, categorized by the virtual world, computational architecture, architectural theory, and so on. Eventually, a total of 187 articles and 2 architectural projects are included in this survey.
Various other surveys further locate this scope, as follows: Category 1 (machine learning  (Kreuzberger et al., 2023; Penney and Chen, 2019) or deep learning (Hatcher and Yu, 2018; Kreuzberger et al., 2023; Alom et al., 2019; Akinosho et al., 2020; Khan et al., 2020; Alom et al., 2019; Sarker, 2021a; Pouyanfar et al., 2018; Abdar et al., 2021)): 3D shape generation (Shi et al., 2023b; Oussidi and Elhassouny, 2018; Cao et al., 2020), scene synthesis  (Xia and Xue, 2023; Zhang et al., 2019), applications (Dong et al., 2021; Dargan et al., 2020), 3D representation (Guo et al., 2021b), 3D reconstruction from 2D  (Yuniarti and Suciati, 2019), and generative models (GM et al., 2020; Aggarwal et al., 2021; Wang et al., 2021; Jabbar et al., 2021; Pavan Kumar and Jayagopal, 2021; Croitoru et al., 2023; Creswell et al., 2018; GM et al., 2020). Category 2 (machine learning in the architecture (Özerol and Selçuk, 2023), or DL-assisted architectural design  (Penney and Chen, 2019; Newton, 2019), or computer-aided architectural design (Stojanovski et al., 2022)): infrastructure  (Tamke et al., 2018), intelligent construction  (Baduge et al., 2022; Akinosho et al., 2020), life cycle  (Hong et al., 2020), and other design  (Regenwetter et al., 2022). And Category 3 (architecture in a virtual environment or virtual worlds): design disciplinary in virtual reality (theories and applications)  (Bartle, 2003), HCI in the virtual architecture (human senses and emotions)  (Lee et al., 2021a; Homolja et al., 2020), metaverse or virtual worlds  (Bartle, 2010; Dionisio et al., 2013; Lee et al., 2021b). In contrast, this article reviews the deep learning approaches in the virtual discipline of architecture regarding virtual environment rules, techniques, design principles, and HCI methods. We advocate for an interdisciplinary approach that integrates these research areas. Our survey article uniquely considers the prominent features of the above categories in this topic and further paves the way for AI-architecture integration in virtual environments with the contributions below.
@@
    We propose research agendas and future directions for generated virtual architecture, considering virtual disciplines, deep learning, and architectural design, thus call for interdisciplinary human-AI collaboration research.

### v4

# Towards AI-Architecture Liberty: A Comprehensive Survey on Design and Generation of Virtual Architecture by Deep Learning

@@ Abstract.
3D shape generation techniques leveraging deep learning have garnered significant interest from both the computer vision and architectural design communities, promising to enrich the content in the virtual environment. However, research on virtual architectural design remains limited, particularly regarding designer-AI collaboration and deep learning-assisted design. In our survey, we reviewed 149 related articles (81.2% of articles published between 2019 and 2023) covering architectural design, 3D shape techniques, and virtual environments. Through scrutinizing the literature, we first identify the principles of virtual architecture and illuminate its current production challenges, including datasets, multimodality, design intuition, and generative frameworks. We then introduce the latest approaches to designing and generating virtual buildings leveraging 3D shape generation and summarize four characteristics of various approaches to virtual architecture. Based on our analysis, we expound on four research agendas, including agency, communication, user consideration, and integrating tools. Additionally, we highlight four important enablers of ubiquitous interaction with immersive systems in deep learning-assisted architectural generation. Our work contributes to fostering understanding between designers and deep learning techniques, broadening access to designer-AI collaboration. We advocate for interdisciplinary efforts to address this timely research topic, facilitating content designing and generation in the virtual environment.
AI-assisted architectural design, 3D geometry generation, designer-AI collaboration, AIGC
@@ 1. Introduction
In recent decades, the exploration of architectural space has shifted from reinforced concrete to digital architecture and further to information frameworks, with virtualization extending beyond the physical layer (Carpo, 2017; Bava, 2020). Architecture is no longer perceived as static, permanent objects but as integral components of a data network and evolving communication between different architectural systems (Claypool, 2019; Schroeder et al., 2001). Digital innovations and technological advancements associated with spline, pixels, voxels, and bits have redefined architectural forms called digital architecture (Jovanovic, 2022). In such context, two powerful forces have driven the birth of generated virtual architecture: AI-assisted design and virtual reality (VR) technique. In recent decades, AI architecture has become a critical sub-genre of architectural discourse advocated by scholars and practitioners (Stojanovski et al., 2022), utilizing deep learning neural networks to concern social issues. Moreover, the burgeoning innovates and techniques of virtual environments (e.g. virtual scenes with virtual reality (VR), augmented reality (AR), immersive videogames, and social VR) have expanded the boundaries of production and design in digital fabrication buildings, such as physical-virtual merging (Cheng et al., 2019; Lee et al., 2021a; Lin and Lo, 2021), bio-signal feedback (Ding et al., 2022a; Pei et al., 2020; Homolja et al., 2020; Pei et al., 2021), and participatory design in the virtual environment (Kim et al., 2022; Kwiecinski et al., 2017). Thus, deep learning facilitates the design and generation of diverse virtual architectures with interactive features, supporting the promising vision of ubiquitous intelligent environments in VR. It simulates real-world scenarios by generating and utilizing 3D building datasets, aligning design objectives with human needs (Özerol and Selçuk, 2023).
@@
Figure 1. Virtual architecture provides rich use affordances, benefiting various aspects, including users, communities, and organizations. (a) Hub Rooms in VRChat; (b) Simulated city scene with various buildings; (c) Interior of individual house in a virtual world game, Second Life; (d) Saint Motel Fan Meet on Mozzila hub; (e) Sci-fi nightclub; (f) Center building in pride month event held on Dencentraland; (g) Realism Sotheby’s virtual gallery; (h) The interactive digital experience in a virtual building for the brand Jose Cuervo designed by Rojkind Arquitectos; (i) The office building of Vice Media Group was designed in metaverse by BIG Architect; (j) ”NFTism” is a virtual NFT gallery developed collaboratively by Zaha Hadid Architects (ZHA) and Journee. The architectural design focuses on user experience, social interaction, and ”dramaturgical” compositions, supporting MMO (massively multiplayer online) technology and integrating audio-video interaction [^2] A report by Archdaily, source: https://www.archdaily.com/972886/zaha-hadid-architects-presents-virtual-gallery-exploring-architecture-nfts-and-the-metaverse; (k) The virtual city designed and developed by ZHA, filled with fluid style virtual architecture; (l) Each piece of land and each item in the virtual land as a non-fungible token could trade in the virtual world of Decentraland; (m) A mapping showing ownership of non-fungible token land in Sandbox; (n) A user-friendly tool for editing and creating buildings and land in social VR, Mozzila hub.
@@
The designers creating virtual buildings with AI include broader users beyond conventional expert designers, encompassing individuals with varying levels of expertise in problem understanding, subject matter, and machine learning (Yang et al., 2018). This innovative designer-AI collaboration, employed by diverse designers, could bring multi-dimensional benefits to the virtual environment. Firstly, designers could utilize AI to create virtual buildings for efficient and large-scale production, thereby expanding virtual environment scenarios with a variety of content (Jang and Kim, 2024). Secondly, designers could prioritize user experiences by considering spatial, social, and physical aspects of human-building interaction (HBI), such as evolving spatial forms and adaptive architectural elements (Alavi et al., 2019). Thirdly, designers could generate virtual buildings that offer social sustainability and economic benefits by embedding computational frameworks or workflows automatically (Jang and Kim, 2024), aligning with the development goals of the virtual environment. These benefits include time and resource savings, as well as fostering continuous renewal and evolutionary capabilities. Additionally, deep learning-generated virtual architecture provides convenient access for both experts and inexperienced users in content production through designer-AI collaboration, contributing to the content ecosystem in the virtual environment. One concrete example is that users could create a house or edit land using a user-friendly tool in social VR (Fig. 2n).
@@
Deep learning’s powerful capability enables the fusion of virtual and physical realities, facilitating diverse lifestyles and a wide range of user, community, and organizational uses (as shown in Fig. 2). Virtual buildings could enhance user presence by immersing themselves in vivid, realistic scenes of the virtual environment (Jang and Kim, 2024) (Fig.  2a-c). Virtual buildings also serve as expansive communication interfaces and containers, fostering social activities and goals in immersive environments (Schroeder et al., 2001). For instance, users require virtual buildings to hold various activities, events, and communities, such as fan meets (Fig. 2d), club parties (Fig. 2e), and Pride Month parades (Fig. 2f). Organizations, including companies, cultural institutions, and charity organizations, also require virtual buildings to maintain urban business premises (Fig. 2g-i). Notably, renowned architectural firms have begun designing virtual cities and buildings in virtual worlds (Fig. 2h-k). Furthermore, the interactive nature of virtual architecture facilitates diverse social systems. It delivers sophisticated information regarding social interactive elements in multiple scenarios, as demonstrated by ZAHA’s advanced virtual architectural projects incorporating interactive technology and architectural design (Fig. 2h-j). Lastly, the virtual architecture enables trading, giveaways, and auctions of virtual land as non-fungible tokens (Fig. 2l-m), enriching the metaverse   [^3] The metaverse refers to a hypothetical, computer-generated world that integrates with reality through various techniques and concepts (Lee et al., 2021b). ecosystem in the future.
@@
The early adoption of emerging technology often encounters accessibility issues, and knowledge from one field may not be easily transferred to another (Davies et al., 2018). In architecture, design research primarily focuses on leveraging neural networks, such as StyleGAN and Pix2Pix, to generate images like floor plans (Chaillou, 2022, 2020), exterior facades (Mohammad et al., 2019; Dong et al., 2023), and stylized images (del Campo et al., 2019; Del Campo, 2021). However, there is a lack of effort in collaborating with deep generative models (DGMs) for direct 3D geometry generation. This oversight results in a deficiency of diversified design philosophies, efficient production logic, and precise geometric control in current deep learning-generated architectural designs. Despite proposals for computational and algorithmic approaches (e.g., (Jang and Kim, 2024; Lab, 2023)), design dimensions of virtual architecture have yet to be systematically explored. This gap impedes the advancement of virtual buildings within virtual environments. Our survey aims to bridge this gap by investigating deep-learning-based 3D shape generation methods for virtual architecture. We emphasize the importance of considering designer-AI collaboration perspectives, particularly inclusivity, to cater to non-tech-savvy architects and layperson users seeking to create virtual buildings in the virtual environment.
@@ 1.2. Methodology and Related Articles
Figure 3. (a) This survey investigates the intersection of various areas. (b) A profile of the number of cited works categorized by different years and topics. The survey’s scope and profile of related articles: a – architectural studies on DGMs; c – computer vision studies; v – works on rules in VWs.
As for the abovementioned problems, the intersection of 3D generation techniques and virtual architecture is still nearly blank. Therefore, we anchored the three fields to conduct this survey to complement the key insights with each other: 3D shape generation techniques, DL-assisted architectural design, and the design considerations in VWs (see Fig. 3(a)). This article collects data using the Preferred Reporting Items for Systematic Reviews and Meta-Analyses (PRISMA) framework (Page et al., 2021).
Identification. We reviewed a sample of 149 articles and primarily focused on works published between 2019 and 2023 (five years, 81.2%) as follows: 2024: 2 (1.3%), 2023: 14 (9.4%), 2022: 25 (16.8%), 2021: 34 (22.8%), 2020: 25 (16.8%), 2019: 23 (15.4%), before 2018: 26 (17.5%) from mentioned three fields (see Fig.  3(b)). We found the articles primarily through publication databases such as ACM Digital Library, IEEE Xplore, ScienceDirect, Springer Link, and CuminCAD. These databases ensure a wide range of computer graphics, virtual environments, and architectural design scope. We determined our search query, including the primary and second terms as searching keywords, after discussion with inner researchers through multiple sessions of iteration and refinements. These keywords were inspired by previous review papers. The primary keywords in the search query are straightforward, reflecting our survey topic, as follows:
@@
    Excluded articles solely discussing real-world architectural design aspects (e.g., natural environment, BIM, construction, floor plans).
@@
    Excluded articles not primarily focused on 3D design generation (e.g., analysis of design cognition, limited 2D generative drawing without 3D transition approaches).
@@
    Excluded articles primarily discussing generative approaches other than deep learning for architecture (e.g., reinforcement learning, genetic algorithms). Only articles discussing deep generative models for building generation were considered.
@@
    Excluded articles discussing broad architecture design topics (e.g., the architecture of algorithms).
@@
    Excluded articles related to 3D generation focusing on the classification and retrieval of computer vision approaches.
When the keywords and abstracts did not appear as the key information or elements within our investigative scope, we read the entire publication to determine its inclusion. After screening, we selected 111 articles and 19 CV research papers published on arXiv, totaling 130 authoritative articles. Additionally, we conducted direct searches through the Google search engine, concluding with 19 articles and 2 relevant architectural projects from the perspectives of virtual worlds, computational architecture, and architectural theory. Ultimately, this survey includes a total of 149 articles and 2 architectural projects.
Various other surveys further locate this scope, as follows: Category 1 (machine learning  (Kreuzberger et al., 2023; Penney and Chen, 2019) or deep learning (Hatcher and Yu, 2018; Kreuzberger et al., 2023; Alom et al., 2019; Akinosho et al., 2020; Khan et al., 2020; Alom et al., 2019; Sarker, 2021a; Pouyanfar et al., 2018; Abdar et al., 2021)): 3D shape generation (Shi et al., 2023b; Oussidi and Elhassouny, 2018; Cao et al., 2020), scene synthesis  (Xia and Xue, 2023; Zhang et al., 2019), applications (Dong et al., 2021; Dargan et al., 2020), 3D representation (Guo et al., 2021b), 3D reconstruction from 2D  (Yuniarti and Suciati, 2019), and generative models (GM et al., 2020; Aggarwal et al., 2021; Wang et al., 2021; Jabbar et al., 2021; Pavan Kumar and Jayagopal, 2021; Croitoru et al., 2023; Creswell et al., 2018; GM et al., 2020); Category 2 (artificial intelligence in the architectural design): machine learning-assisted architectural design (Özerol and Selçuk, 2023; Penney and Chen, 2019; Tamke et al., 2018), GAN-assisted architectural design (Newton, 2019), computer-aided architectural design (Stojanovski et al., 2022), historical development (Chaillou, 2022), infrastructure  (Tamke et al., 2018), intelligent construction  (Baduge et al., 2022; Akinosho et al., 2020; Hong et al., 2020); and Category 3 (architecture in a virtual environment or virtual worlds): design disciplinary in virtual reality (theories and applications)  (Bartle, 2003), human perspective in the virtual space  (Lee et al., 2021a; Homolja et al., 2020), human-building interaction (HBI) (Alavi et al., 2019), and metaverse or virtual worlds  (Bartle, 2010; Dionisio et al., 2013; Lee et al., 2021b). In contrast, this article reviews the deep learning approaches in the virtual discipline of architecture regarding virtual environment rules, techniques, design principles, and HCI methods. We advocate for an interdisciplinary approach that integrates these research areas. Our survey article uniquely considers the prominent features of the above categories in this topic and further paves the way for AI-architecture integration in virtual environments with the contributions below.
@@
    We propose research agendas and future directions for generated virtual architecture, considering virtual disciplines, deep learning, and architectural design, thus call for interdisciplinary designer-AI collaboration research.